# Exact Coset Sampling forQuantum Lattice Algorithms

### v6

@@ Abstract
We revisit the post-processing phase of Chen’s Karst-wave quantum lattice algorithm (chen2024quantum) in the Learning with Errors (LWE) parameter regime. Conditioned on a transcript $E$, the post-Step 7 coordinate state on $(\mathbb{Z}_{M})^{n}$ is supported on the shifted grid line
@@
where $\Delta=2D^{2}\mathbf{b}$ and $M=2M_{2}=2D^{2}Q$ with $Q$ odd. The amplitudes contain a quadratic chirp $\omega_{Q}^{-j^{2}}$ and an unknown run-dependent offset $\bm{v}^{\ast}(E)$. We show that Chen’s Steps 8-9 can be replaced by a single routine, Step $9^{\dagger}$: measure $\tau:=X_{1}\bmod D^{2}$, apply a $\tau$-dependent quadratic phase on the first coordinate register to remove the chirp, and then apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the coordinate registers. The chirp removal uses a canonical gauge promise: conditioned on $E$, the offset $v_{1}^{\ast}(E)\bmod M_{2}$ equals the centered lift $\operatorname{ctr}(\tau)$. Under explicit front-end conditions (AC1-AC5), a Fourier sample $\bm{u}\in\mathbb{Z}_{M}^{n}$ satisfies the resonance $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$ with probability $1-\operatorname{negl}(n)$, and conditioned on resonance the reduced sample $\bm{u}\bmod Q$ is exactly uniform on the dual hyperplane. This yields exact recovery of $\mathbf{b}\bmod Q$ and hence the embedded LWE secret.
@@ 1 Introduction
for an outcome-dependent amplitude profile $\alpha_{E}(j,\bm{k})$ and an index set $\mathcal{K}\subseteq\{0\}\times\mathbb{Z}^{n-1}$ that comes from the grid-state construction. When it is clear from context, we suppress the transcript index $E$. The phase factor $i^{\|\bm{k}\|^{2}}$ comes from Chen’s grid-state preparation procedure. We view the coordinate values as elements of $(\mathbb{Z}_{M})^{n}$ throughout. Accordingly, $\mathbf{X}(j):=2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}$ is interpreted componentwise modulo $M$ (as in Chen’s post-Step 7 expression), and $\bm{k}$ records the $M_{2}$-shift term $M_{2}\bm{k}$. Reducing the coordinate registers modulo $M_{2}$ eliminates this shift and yields the affine congruence $\mathbf{X}(j)\equiv 2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}\pmod{M_{2}}$ used in Additional Condition AC1.
@@
Additional Condition AC1 records the affine relation $\mathbf{X}(j)=2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}$ for the post-Step 7 state. Additional Condition AC4 fixes a canonical gauge for the first-coordinate offset: the preparation circuit is compiled (by an $E$-controlled basis translation on $X_{1}$ by a multiple of $D^{2}$) so that measuring $\tau:=X_{1}\bmod D^{2}$ determines $v_{1}^{\ast}(E)\bmod Q$ for that run. Under Additional Condition AC1, the first coordinate satisfies the precise congruence
@@
where under the canonical gauge (AC4) the post-processing computes $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$ from $\tau=X_{1}\bmod D^{2}$.
In Chen’s post-processing one ultimately outputs non-zero Fourier samples in $\mathbb{Z}_{M_{2}}^{n}$ (with $M_{2}=M/2$) whose reduction modulo $Q=M/(2D^{2})$ satisfies a homogeneous dual relation. In our direct-QFT variant we instead measure $\bm{u}\in\mathbb{Z}_{M}^{n}$ and store only the reduction $\bm{u}_{Q}:=\bm{u}\bmod Q$. In either view, the goal is to obtain non-zero vectors satisfying
@@
where $Q:=\frac{M}{2D^{2}}$ is an odd integer. (In chen2024quantum the parameters are additionally arranged so that $Q$ has a convenient factorization, but Step $9^{\dagger}$ does not use it.) The short vector $\mathbf{b}=(b_{1},\dots,b_{n})\in\mathbb{Z}^{n}$ stores the hidden data and satisfies the normalization $b_{1}=-1$. The offset $\bm{v}^{\ast}$ is unknown and can change from run to run. The algorithm then collects $O(n)$ such samples and solves the resulting linear system modulo $Q$ to recover $\mathbf{b}\bmod Q$ and hence the hidden data vector.
@@ 2 Background and Access Model
When we write $\langle\bm{a},\bm{u}\rangle\equiv 0\pmod{Q}$ this is the intended meaning (i.e., the dot product after coordinate-wise reduction $\mathbb{Z}_{M}\to\mathbb{Z}_{Q}$).
@@
For clarity we treat $\mathrm{QFT}_{\mathbb{Z}_{M}}$ and the controlled root-of-unity phases used in the chirp correction (AC4) as exact unitaries. In a standard uniform gate set these operations can be implemented to precision $2^{-\operatorname{poly}(n)}$, which perturbs all output distributions by at most $\operatorname{negl}(n)$ in total variation distance. We suppress this approximation issue throughout, since it does not affect any of the modular-algebraic arguments.
@@
In Chen’s notation, $M=2(t^{2}{+}u^{2})$ and $x=D\mathbf{b}$, where $\mathbf{b}$ is the short vector from the LWE-to-lattice reduction (Eq. (12) in chen2024quantum). We follow this identification. Chen’s Step 9 defines a relabelled vector $\mathbf{b}^{\ast}$ (Eq. (39) in that paper). Our post-processing never uses $\mathbf{b}^{\ast}$. We keep all formulas in terms of $\mathbf{b}$ instead.
Under Chen’s parameter constraints the quantity $M/(2D^{2})$ is an odd integer. We write
@@
In Chen’s concrete instantiation $Q$ happens to admit a factorisation into pairwise coprime odd factors; the new Step $9^{\dagger}$ does not use this extra structure and only relies on $Q$ being odd.
@@
The scaling parameter $D$ can then be chosen (odd, $\gcd(D,Q)=1$, and $\operatorname{poly}(n)$-bounded) to satisfy the geometric separation requirement of the Karst-wave front end (Condition C.6 of chen2024quantum). The only additional analytic requirement used by Step $9^{\dagger}$ is the one-dimensional spectral concentration property recorded in Additional Condition AC5, which is achieved in Chen’s complex-Gaussian regime once the loop-envelope width satisfies $\sigma_{J}\gtrsim Q\log n$ (Remark 2.5). All subsequent correctness statements are conditional on the existence of a (re-tuned) front-end configuration satisfying Definition 2.1 and Additional Conditions AC1-AC5.
@@
Take LWE parameters $(\ell,m,q,\beta)$ in the regime of chen2024quantum and set $n:=1+\ell+m$. Let $L_{q}^{\perp}(\mathbf{A})$, the planted shortest vector $\mathbf{b}$, and the integers $(p_{\eta})_{\eta}$ be as in Section 3.2 of chen2024quantum. A re-tuned Karst-wave configuration is any choice of front-end parameters
@@
    Choose $r,s>0$ so that Chen’s Karst-wave constraints C.5-C.7 hold and the resulting geometric width parameter
@@
    For the resulting complex-Gaussian front end, let $\sigma_{J}$ denote the standard deviation of the one-dimensional envelope $\alpha_{E}(j)$ along the loop index $j$ (after marginalizing over the grid index and the workspace and conditioning on a transcript $E$). We restrict to parameter choices satisfying
@@
    for some fixed absolute constant $C_{\mathrm{spec}}>0$. A lower bound of this form keeps the spectral leakage in Additional Condition AC5 bounded by $\exp(-\Omega(\log^{2}n))$.
@@ 2.1 Front-end properties and additional conditions
for some index set $\mathcal{K}\subseteq\{0\}\times\mathbb{Z}^{n-1}$, where $\Delta=2D^{2}\mathbf{b}$ and the offset $\bm{v}^{\ast}(E)\in(\mathbb{Z}_{M})^{n}$ depends on $E$ but not on $j$. Reducing modulo $M_{2}$ eliminates the $M_{2}\bm{k}$ term and yields the affine congruence
@@
We assume throughout that $Q$ is an odd integer. The post-processing in Step $9^{\dagger}$ works for an arbitrary composite modulus $Q$; whenever we invert a residue class we only do so modulo a modulus that is guaranteed to be coprime to that residue (in particular, we only invert $2D^{2}$ modulo $Q$, which is valid by AC2-AC3). In particular, we do not require $Q$ to factor in any prescribed way, and the integers $(p_{\eta})_{\eta}$ from the LWE-to-lattice reduction play no role in the algebraic structure of $Q$.
@@
The post-Step 7 amplitudes contain the quadratic chirp $\omega_{Q}^{-j^{2}}$ in the loop index $j$. We cancel it with a diagonal unitary acting only on the first coordinate register $X_{1}$.
Fix a transcript $E$. By AC1 and $b_{1}=-1$, for every $j$ in the effective window,
$X_{1}\equiv-2D^{2}j+v_{1}^{\ast}(E)\pmod{M_{2}}$.
Because $k_{1}=0$ for all $\bm{k}\in\mathcal{K}$, reducing modulo $D^{2}$ removes the $j$-term on every branch:
@@
Thus $\tau$ is deterministic conditioned on $E$; measuring it does not disturb the coherent superposition over $j$ and $\bm{k}$. Let $\bar{v}_{1}:=\operatorname{ctr}(\tau)\in(-D^{2}/2,D^{2}/2]$ and set $v_{1,Q}:=\bar{v}_{1}\bmod Q\in\mathbb{Z}_{Q}$.
Under the canonical-gauge requirement below, this $v_{1,Q}$ equals $v_{1}^{\ast}(E)\bmod Q$ for that run.
Canonical gauge (required). Measuring $\tau$ determines $v_{1}^{\ast}(E)\bmod D^{2}$ but leaves an ambiguity by multiples of $D^{2}$ inside $\mathbb{Z}_{M_{2}}$. Equivalently, there is a unique $\lambda(E)\in\mathbb{Z}_{Q}$ such that
$v_{1}^{\ast}(E)\equiv\bar{v}_{1}+\lambda(E)D^{2}\pmod{M_{2}}$.
Any correction depending only on $\tau$ cancels $\omega_{Q}^{-j^{2}}$ only up to a residual run-dependent linear phase $\omega_{Q}^{-\lambda(E)j}$ (up to a global constant), which shifts the resonance class $\langle\mathbf{b},\bm{u}\rangle\bmod Q$ and prevents aggregating samples across invocations. We therefore assume a canonical gauge in which $\lambda(E)=0$ for every transcript $E$, i.e.,
$v_{1}^{\ast}(E)\equiv\bar{v}_{1}\pmod{M_{2}}$.
When the preparation circuit is known, this can be enforced by an $E$-controlled basis translation on $X_{1}$ by a multiple of $D^{2}$, provided $E$ allows reversible computation of $\lambda(E)$; see Proposition 2.3.
Define the classical inverse $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$, which exists by AC2-AC3. For each basis value $x\in\mathbb{Z}_{M}$ of $X_{1}$, write $x_{Q}:=x\bmod Q$ and define
$j_{x}:=\mathrm{inv}\cdot\bigl((x_{Q}-v_{1,Q})\bmod Q\bigr)\in\mathbb{Z}_{Q}.$
For a branch with loop label $j$ we have $(x_{Q}-v_{1,Q})\equiv-2D^{2}j\pmod{Q}$, hence $j_{x}\equiv j\pmod{Q}$. We apply the diagonal unitary $U_{\mathrm{corr}}^{(\tau)}$ on $X_{1}$:
@@
We view the loop index $j\in\mathbb{Z}_{P}$ as embedded in $\mathbb{Z}$. We extend $\alpha_{E}(j,\bm{k})$ by zero outside an interval $\{|j|\leq J_{\max}\}$. The effective window satisfies $J_{\max}=\operatorname{poly}(Q\log n)\ll P$ in the Karst-wave regime below; in particular we assume $J_{\max}<P/2$ so that the embedding of $j\in\mathbb{Z}_{P}$ into integers is unambiguous on the support of the state. The infinite sum above then matches, up to negligible error, the finite sum that appears in Chen’s state.
As in Chen’s Karst-wave front end, we assume that conditioned on $E$ the amplitudes factor across the loop index and the grid label:
@@
Additional Conditions AC2-AC3 are enforced directly by Definition 2.1. Additional Condition AC1 is the standard offset-coherence invariant of Chen’s deferred-measurement front end (Lemma 2.8). Additional Condition AC4 is the only genuinely new access-model promise used by Step $9^{\dagger}$: it replaces Chen’s Step 8 offset-extraction gadget by a gauge-fixing guarantee (or coherent access to the gauge parameter) sufficient to cancel the quadratic chirp using only $X_{1}$. It is enforceable by a compile-time basis translation on the first coordinate whenever the preparation circuit supports reversible computation of the gauge parameter $\lambda(E)$ (Proposition 2.3). Finally, Additional Condition AC5 follows from the complex discrete-Gaussian Karst-wave envelope once $\sigma_{J}\geq C_{\mathrm{spec}}Q\log n$ (Proposition 2.6).
@@
Conditioned on $R$, the reduced outcome $\bm{u}\bmod Q\in\mathbb{Z}_{Q}^{n}$ is exactly uniform on $H$. In particular, $\Pr[\bm{u}\bmod Q=\bm{0}\mid R]=1/|H|=Q^{-(n-1)}$.
@@
Fix $E$. Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the chirp-free state (1.2). For each $\bm{u}\in\mathbb{Z}_{M}^{n}$, the (unnormalized) amplitude equals

### v7

@@ Abstract
We revisit the post-processing phase of Chen’s Karst-wave quantum lattice algorithm (chen2024quantum) in the Learning with Errors (LWE) parameter regime. Conditioned on a transcript $E$, the post-Step 7 coordinate state on $(\mathbb{Z}_{M})^{n}$ is supported on an affine grid line
@@
with $\Delta=2D^{2}\mathbf{b}$, $M=2M_{2}=2D^{2}Q$, and $Q$ odd. The amplitudes include a quadratic Karst-wave chirp $\omega_{Q}^{-j^{2}}$ and an unknown run-dependent offset $\bm{v}^{\ast}(E)$. We show that Chen’s Steps 8-9 can be replaced by a single exact post-processing routine: measure the deterministic residue $\tau:=X_{1}\bmod D^{2}$, obtain the run-local class $v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q$ as explicit side information in our access model, apply a $v_{1,Q}$-dependent diagonal quadratic phase on $X_{1}$ to cancel the chirp, and then apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the coordinate registers. The routine never needs the full offset $\bm{v}^{\ast}(E)$. Under Additional Conditions AC1-AC5 on the front end, a measured Fourier outcome $\bm{u}\in\mathbb{Z}_{M}^{n}$ satisfies the resonance $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$ with probability $1-\operatorname{negl}(n)$. Moreover, conditioned on resonance, the reduced outcome $\bm{u}\bmod Q$ is exactly uniform on the dual hyperplane $H=\{\bm{v}\in\mathbb{Z}_{Q}^{n}:\langle\mathbf{b},\bm{v}\rangle\equiv 0\pmod{Q}\}$.
@@ 1 Introduction
for an outcome-dependent amplitude profile $\alpha_{E}(j,\bm{k})$ and an index set $\mathcal{K}\subseteq\{0\}\times\{0,1\}^{n-1}$ that comes from the grid-state construction. When it is clear from context, we suppress the transcript index $E$. The phase factor $i^{\|\bm{k}\|^{2}}$ comes from Chen’s grid-state preparation procedure. We view the coordinate values as elements of $(\mathbb{Z}_{M})^{n}$ throughout. Accordingly, $\mathbf{X}(j):=2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}$ is interpreted componentwise modulo $M$, and $\bm{k}$ records the $M_{2}$-shift term $M_{2}\bm{k}$. Reducing the coordinate registers modulo $M_{2}$ eliminates this shift and yields the affine congruence $\mathbf{X}(j)\equiv 2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}\pmod{M_{2}}$ used in Additional Condition AC1.
@@
Additional Condition AC4 isolates the only extra access needed to cancel the Karst-wave chirp using only the first coordinate register: after measuring the deterministic residue $\tau:=X_{1}\bmod D^{2}$, the post-processing must obtain the run-local class $v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q$. This information is not implied by AC1 alone; in our access model, it is provided as a short run-local gauge value. If the gauge value is available coherently inside the preparation, one can compile a canonical-gauge variant with $\lambda(E)=0$ (Proposition 2.3), in which case $v_{1,Q}=\operatorname{ctr}(\tau)\bmod Q$. Under Additional Condition AC1, the first coordinate satisfies the precise congruence
@@
where under AC4 the post-processing computes $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$ from $\tau=X_{1}\bmod D^{2}$ together with the run-local gauge value.
In Chen’s post-processing, one ultimately outputs non-zero Fourier samples in $\mathbb{Z}_{M_{2}}^{n}$ (with $M_{2}=M/2$) whose reduction modulo $Q=M/(2D^{2})$ satisfies a homogeneous dual relation. In our direct-QFT variant, we instead measure $\bm{u}\in\mathbb{Z}_{M}^{n}$ and store only the reduction $\bm{u}_{Q}:=\bm{u}\bmod Q$. In either view, the goal is to obtain non-zero vectors satisfying
@@
where $Q:=\frac{M}{2D^{2}}$ is an odd integer. The short vector $\mathbf{b}=(b_{1},\dots,b_{n})\in\mathbb{Z}^{n}$ stores the hidden data and satisfies the normalization $b_{1}=-1$. The offset $\bm{v}^{\ast}$ is unknown and can change from run to run. The algorithm then collects $O(n)$ such samples and solves the resulting linear system modulo $Q$ to recover $\mathbf{b}\bmod Q$ and hence the hidden data vector.
@@ 2 Background and Access Model
When we write $\langle\bm{a},\bm{u}\rangle\equiv 0\pmod{Q}$, this is the intended meaning.
@@
In the intended LWE instantiation (Definition 2.1) we have $Q=\operatorname{poly}(n)$, hence $Q^{-(n-1)}=\operatorname{negl}(n)$.
@@
For clarity, we treat $\mathrm{QFT}_{\mathbb{Z}_{M}}$ and the controlled root-of-unity phases used in the chirp correction (AC4) as exact unitaries. In a standard uniform gate set, these operations can be implemented to precision $2^{-\operatorname{poly}(n)}$, which perturbs all output distributions by at most $\operatorname{negl}(n)$ in total variation distance. We suppress this approximation issue throughout, since it does not affect any of the modular-algebraic arguments. See, e.g., standard treatments of approximate QFT and fault-tolerant synthesis (nielsen2010quantum).
@@
In Chen’s notation, $M=2(t^{2}{+}u^{2})$ and $x=D\mathbf{b}$, where $\mathbf{b}$ is the short vector from the LWE-to-lattice reduction (Eq. (12) in chen2024quantum). We follow this identification. Chen’s Step 9 defines a relabelled vector $\mathbf{b}^{\ast}$. Our post-processing never uses $\mathbf{b}^{\ast}$. We keep all formulas in terms of $\mathbf{b}$ instead.
Under Chen’s parameter constraints, the quantity $M/(2D^{2})$ is an odd integer. We write
@@
In Chen’s concrete instantiation, $Q$ happens to admit a factorization into pairwise coprime odd factors, the new Step $9^{\dagger}$ does not use this extra structure and only relies on $Q$ being odd.
@@
The scaling parameter $D$ can then be chosen (odd, $\gcd(D,Q)=1$, and $\operatorname{poly}(n)$-bounded) to satisfy the geometric separation requirement of the Karst-wave front end (Condition C.6 of chen2024quantum). The only additional analytic requirement used by Step $9^{\dagger}$ is the one-dimensional spectral concentration property recorded in Additional Condition AC5, which is achieved in Chen’s complex-Gaussian regime once the loop-envelope width satisfies $\sigma_{J}\gtrsim Q\log n$ (Remark 2.5). All subsequent correctness statements are conditional on the existence of a re-tuned front-end configuration satisfying Definition 2.1 and Additional Conditions AC1-AC5.
@@
Take LWE parameters $(\ell,m,q,\beta)$ in the regime of chen2024quantum and set $n:=1+\ell+m$. Let $L_{q}^{\perp}(\mathbf{A})$, the planted shortest vector $\mathbf{b}$, and the integers $(p_{\eta})_{\eta}=(p_{1},\dots,p_{\kappa})$ be as in Section 3.2 of chen2024quantum. We assume $\kappa\geq 2$ and $\kappa-1\leq\ell$, so that the $(\kappa-1)$ chosen secret coordinates fit in the $\ell$-dimensional secret. A re-tuned Karst-wave configuration is any choice of front-end parameters
@@
    Choose $r,s>0$ so that Chen’s Karst-wave constraints C.5-C.7 hold, and the resulting geometric width parameter
@@
    For the resulting complex-Gaussian front end, let $\sigma_{J}$ denote the standard deviation of the one-dimensional envelope $\alpha_{E}(j)$ along the loop index $j$. We restrict to parameter choices satisfying
@@
    for some fixed absolute constant $C_{\mathrm{spec}}>0$. A lower bound of this form keeps the spectral leakage in Additional Condition AC5 bounded by $\exp(-\Omega(\log^{2}n))$. In contrast, the geometric width parameter $\sigma_{\mathrm{geom}}$ in (R3) controls the within-ball width of each extracted Gaussian ball in coordinate space; these are distinct quantities in Chen’s regime.
@@ 2.1 Front-end properties and additional conditions
for some index set $\mathcal{K}\subseteq 0\!\mid\!\{0,1\}^{n-1}$, where $\Delta=2D^{2}\mathbf{b}$ and the offset $\bm{v}^{\ast}(E)\in(\mathbb{Z}_{M})^{n}$ depends on $E$ but not on $j$. As above, $\mathcal{K}$ is a choice of parity representatives; any larger grid index domain can be reduced to this form by regrouping all terms with the same $\bm{k}\bmod 2$ into a single amplitude. Reducing modulo $M_{2}$ eliminates the $M_{2}\bm{k}$ term and yields the affine congruence
@@
We assume throughout that $Q$ is an odd integer. The post-processing in Step $9^{\dagger}$ works for an arbitrary composite modulus $Q$. Whenever we invert a residue class, we only do so modulo a modulus that is guaranteed to be coprime to that residue. In particular, we do not require $Q$ to factor in any prescribed way, and the integers $(p_{\eta})_{\eta}$ from the LWE-to-lattice reduction play no role in the algebraic structure of $Q$.
@@
Run-local chirp cancellation from the first coordinate. The post-Step 7 amplitudes contain the quadratic chirp $\omega_{Q}^{-j^{2}}$ in the loop index $j$. Fix a transcript $E$. By AC1 and $b_{1}=-1$, for every $j$ in the effective window,
$X_{1}\equiv-2D^{2}j+v_{1}^{\ast}(E)\pmod{M_{2}},\qquad M_{2}=D^{2}Q.$
Since $k_{1}=0$ for all $\bm{k}\in\mathcal{K}$, reducing modulo $D^{2}$ removes the $j$-term on every branch:
@@
Thus $\tau$ is deterministic conditioned on $E$; measuring it does not disturb the coherent superposition over $j$ and $\bm{k}$. Let $\bar{v}_{1}:=\operatorname{ctr}(\tau)\in(-D^{2}/2,D^{2}/2]$.
Side information. We assume the preparation oracle also returns, as classical side information, either the residue
$v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q\in\mathbb{Z}_{Q},$
or equivalently a “high-order” gauge value $\lambda(E)\in\mathbb{Z}_{Q}$ satisfying $v_{1}^{\ast}(E)\equiv\bar{v}_{1}+\lambda(E)D^{2}\pmod{M_{2}}$, so that $v_{1,Q}=(\bar{v}_{1}+\lambda(E)D^{2})\bmod Q$. (See Proposition 2.3 for a canonical-gauge compilation where $\lambda(E)=0$.)
Chirp-cancellation unitary. Define $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$, which exists by AC2-AC3. For each basis value $x\in\mathbb{Z}_{M}$ of $X_{1}$, write $x_{Q}:=x\bmod Q$ and define
[…]
We apply the diagonal unitary $U_{\mathrm{corr}}^{(v_{1,Q})}$ on $X_{1}$,
@@
We view the loop index $j\in\mathbb{Z}_{P}$ as embedded in $\mathbb{Z}$. We extend $\alpha_{E}(j,\bm{k})$ by zero outside an interval $\{|j|\leq J_{\max}\}$. The effective window satisfies $J_{\max}=\operatorname{poly}(Q\log n)\ll P$ in the Karst-wave regime below; in particular, we assume $J_{\max}<P/2$ so that the embedding of $j\in\mathbb{Z}_{P}$ into integers is unambiguous on the support of the state. The infinite sum above then matches, up to negligible error, the finite sum that appears in Chen’s state.
As in Chen’s Karst-wave front end, we assume that conditioned on $E$, the amplitudes factor across the loop index and the grid label:
@@
Additional Conditions AC2-AC3 are enforced directly by Definition 2.1. Additional Condition AC1 is the standard offset-coherence invariant of Chen’s deferred-measurement front end (Lemma 2.8). Additional Condition AC4 is the only genuinely new access-model requirement used by Step $9^{\dagger}$: it asks for run-local access to $v_{1,Q}(E)$ (equivalently, to a gauge value $\lambda(E)$) so that the quadratic chirp can be cancelled using only $X_{1}$. This information is not implied by AC1 and is treated as explicit side information returned by the oracle; if the gauge value is available coherently inside the preparation, one can also compile a canonical gauge via an $E$-controlled basis translation on $X_{1}$ (Proposition 2.3). Finally, Additional Condition AC5 follows from the complex discrete-Gaussian Karst-wave envelope once $\sigma_{J}\geq C_{\mathrm{spec}}Q\log n$ (Proposition 2.6).
@@
If $\Pr_{\mu_{E}}[R]>0$, then conditioned on $R$ the reduced outcome $\bm{u}\bmod Q\in\mathbb{Z}_{Q}^{n}$ is exactly uniform on $H$. In particular, $\Pr[\bm{u}\bmod Q=\bm{0}\mid R]=1/|H|=Q^{-(n-1)}$.
@@
Fix $E$. Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the chirp-free state (1.2). For each $\bm{u}\in\mathbb{Z}_{M}^{n}$, the unnormalized amplitude equals